OpenStudy (anonymous):

a bottle of soft drink is two thirds full. one quarter of the drink is consumed. how full is the bottle now?

OpenStudy (anonymous):

2/3-1/4 = is the problem first match the denominators in the fractions ... they both share 12 ..so its 8 /12 and 3/12 8/12- 3/12 = 5/12